Exercise 6-8 Compute the Margin of Safety [LO6-7]

Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month’s budget appear below:

Selling price per unit $ 25
Variable expense per unit $ 18
Fixed expense per month $ 5,670
Unit sales per month 960

Required:

1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as a percentage of its sales? (Round your percentage answer to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)
Answer #1

1. The company’s margin of safety = $3,750

2. The company’s margin of safety as a percentage of its sales = 15.63%

Explanation

Contribution per unit = Selling price per unit - Variable expense per unit

= 25 - 18 = $7 per unit

Contribution = 960 units * $7 per unit = $6,720

Contribution margin ratio = Contribution per unit / Selling price per unit

= 7 / 25 = 28%

Profit = Contribution - Fixed expense

= $6,720 - $5,670 = $1,050

Therefore, margin of safety = Profit / Contribution margin ratio

= $1,050 / 28% = $3,750

Therefore, margin of safety as a percentage of its sales = Margin of safety / Sales

Sales = 960 units * $25 per unit = $24,000

Therefore, margin of safety as a percentage of its sales = $3,750 / $24,000 = 15.63%